OpenStudy (anonymous):

find the diameter and radius of a circle with the given circumference. Round to the nearest hundredth 1.) C=36m

OpenStudy (jdoe0001):

\(\bf circumference = 2\pi r\qquad r=radius = \frac{diameter}{2}\implies 2r = diameter\\ \quad \\ 36= 2\pi r\) solve for "r"

OpenStudy (anonymous):

radius is always half the diameter and diameter is always twice the radius

OpenStudy (anonymous):

Circumference is always the same as the diameter so...

OpenStudy (anonymous):

@lalalovekimmy Make sense?

OpenStudy (anonymous):

@jdoe0001 @ShaTay yes :) thank you!

OpenStudy (anonymous):

@lalaloveskimmy of course no problem